State the number of solutions for the following quadratic system?
y^2=x^2-7
x^2+y^2=25

Answer by ewatrrr(24785)

 
Hi,
y^2=x^2-7
y^2= -x^2 + 25
x^2 - 7 = -x^2 + 25
2x^2 = 32
x^2 = 16
x = ± 4